40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 17 16:44:04 2015, Koji, Update, LSC, Delay line installed again 
    Reply  Tue Feb 17 19:49:51 2015, Koji, Update, LSC, Delay line un-installed again 
       Reply  Mon Feb 23 21:57:13 2015, Koji, Update, LSC, Delay line installed again (experiment, round 1) 
          Reply  Tue Feb 24 01:57:01 2015, Jenne, Update, LSC, Modelled effect of relative modulation phase ChangeInAngle_PRCLvsMICH.pngREFL33_MICH_SensMat_vsModulationOrder_vsModulationPhase.pngREFL33_SensMat_vsModulationPhase.pngREFL33_PRCL_nominalVsLoweredGamma.pngREFL33_MICH_nominalVsLoweredGamma.png
             Reply  Thu Feb 26 00:20:54 2015, ericq, Update, LSC, Modelled effect of relative modulation phase 33budget_canceled.png33budget_nominal.png2015-02-ModPhase.zip
                Reply  Thu Feb 26 01:51:39 2015, ericq, Update, LSC, Sideband HOMs C1_HOMcurves_PR.pngC1_HOMcurves_Y.pngC1_HOMcurves_X.pngC1_HOMlist.zip
                Reply  Thu Feb 26 01:53:35 2015, Koji, Update, LSC, Modelled effect of relative modulation phase 
       Reply  Tue Feb 24 23:39:16 2015, Jenne, Update, LSC, Delay line re-installed again 
          Reply  Wed Feb 25 04:59:47 2015, Jenne, Update, LSC, Delay line re-installed, measurements round 2 SensMat_PRMI_24Feb2015.pdf
             Reply  Wed Feb 25 12:16:27 2015, Koji, Update, LSC, Delay line re-installed, measurements round 2 
                Reply  Wed Feb 25 14:18:28 2015, rana, Update, LSC, Delay line re-installed, measurements round 2 
          Reply  Wed Feb 25 14:51:13 2015, Jenne, Update, LSC, Delay line un-installed again 
             Reply  Wed Feb 25 20:00:39 2015, Jenne, Update, LSC, Delay line un-installed again - sensing matrix comparisons 
                Reply  Wed Feb 25 23:48:57 2015, rana, Update, LSC, some thoughts 
                   Reply  Thu Feb 26 20:02:46 2015, Jenne, Update, LSC, re: some thoughts 
Message ID: 11071     Entry time: Wed Feb 25 23:48:57 2015     In reply to: 11070     Reply to this: 11080
Author: rana 
Type: Update 
Category: LSC 
Subject: some thoughts 
  • Comparing just the 2 cases with locking on 33, it seems that the 55 MHz gain has changed by 14 dB instead of the 10 dB that we expected. Is it that we need to measure the modulation drive change more carefully, or just that the PRMI was aligned differently?
  • The 165 signal changed by a factor of 60 (35 dB) which is more consistent with a ~12 dB change in Gamma2, so not so far off.
  • The fact that the whole sensing matrix increases in amplitude between 33 and 55 lock makes me think that either the alignment was very bad for the 33 lock, or that the 33 signals have a significant offset; if that's the case, then we should do as LLO and set the digital offsets in the 33 signals by locking first on 55.
  • How does the REFL33 demod phase change by 70 deg?
ELOG V3.1.3-